# Radio Morphing: Fast computation of inclined air shower radio emission

### v1

@@ Abstract
The preparation of next-generation large-scale radio experiments requires running a fast and efficient number of simulations to explore multiple detector configurations over vast areas and develop novel methods for the reconstruction of air shower parameters. While Monte Carlo simulations are accurate and reliable tools, they are too computationally expensive to explore the full parameter space of these new detectors within a reasonable timescale. We introduce a new version of Radio Morphing, a semi-analytical tool designed to simulate the radio emission of any cosmic-ray induced air shower with zenith angle $\theta>60^{\circ}$, at any desired antenna position, from the simulation data of a few reference showers at given positions. We present the latest performances of Radio Morphing which now provides simulation of air shower radio signals with average relative differences on the peak amplitude below $17\%$ on raw traces, below $15\%$ with a $3\sigma$ trigger threshold, below $13\%$ in the $50-200]\,\rm MHz$ band, and even below $\sim 10\%$ in the $30-80]\,\rm MHz$ band. These results are combined with a computation time reduced by more than four orders of magnitude, compared to standard Monte Carlo simulations.
@@ 1 Introduction
Radio detection is now a mature, well-established technique, used by a wide range of experiments for air shower detection to unveil the high-energy Universe. Large-scale radio experiments such as GRAND [1], BEACON [2], SKA [3], or the radio upgrade of the Pierre Auger Observatory [4] are now being built to reach unprecedented sensitivity, field of view, and angular resolution at the highest energies. This goal requires to run a large number of air-shower simulations to explore different antenna configurations, test reconstruction methods, and evaluate the detector performances. While Monte Carlo simulations such as ZHAireS [5, 6] and Coreas [7, 8] provide accurate simulation results, these are too computationally intensive to explore all the parameter space with large-scale experiments. On the other hand, macroscopic approaches such as MGMR [9, 10] are much faster but limited by our incomplete knowledge of the physical processes, particularly when looking for showers with very inclined arrival directions. This motivates the need to develop fast and accurate simulations of the radio signal. Already significant efforts have been made towards this direction, with template synthesis for vertical and mildly inclined showers [11]. On the other hand, inclined showers remain challenging because of asymmetry effects and rapid changes in the emission with the shower zenith angle. We introduce here a new version of Radio Morphing, a semi-analytical tool designed to simulate the radio emission from any air shower with zenith angle above $60^{\circ}$, at any position, from the reference data of a few simulated showers at given positions. The current version incorporates shower-to-shower fluctuations, refined scaling laws, interpolation method, and reference library compared to the version of [12]. We present the current Radio Morphing method and its implementation in Sections 2 and 3 respectively, and show the latest Radio Morphing performances in Section 4.
@@ 2.1 First principles of Radio Morphing
We mention that the whole Radio Morphing procedure is performed on the full-band electric field traces, to infer general results that can then be filtered in the desired frequency range, similarly to Monte Carlo simulations.
@@ 2.2 Reference library and plane selection
When generating a Radio Morphing simulation, the user must specify a target primary energy, zenith and azimuth angle $(E_{t},\,\theta_{t},\,\varphi_{t})$. We thus chose, in our library, the reference ZHAireS simulation with the closest zenith angle $\theta_{\rm ref}$. The user must also specify a coordinate list of target antenna positions ($x_{t}$, $y_{t}$, $z_{t}$) in the AIRES reference frame, at which the radio signal will be computed. Hence, for each target antenna, we chose within the selected ZHAireS simulation, the reference plane with $X_{\rm max}$ distance $D_{\rm plane}$ that is the closest to the $X_{\rm max}$-antenna distance $d_{\rm ant}$. This procedure is done to reduce as much as possible the error related to the extrapolation of the signal.
@@ 3.1 Primary composition
Figure 2: (Top) Lateral distribution functions (LDFs) for two showers with the same arrival directions ($\theta=75^{\circ},\,\varphi=90^{\circ}$), but different primary energies ($E_{r}=3.98\,\rm EeV,\,E_{t}=1.0\,\rm EeV$) without any energy scaling (top left) and with the energy scaling procedure (top right). The $\omega$ angle corresponds to the angular deviation from the shower axis, taken from $X_{\rm max}$ as shown in Fig. 1. (Bottom) LDFs for two showers with the same primary energy and zenith angle ($E=3.98\,\rm EeV,\theta=75^{\circ}$), but different azimuth angles ($\varphi_{r}=90^{\circ},\,\varphi_{t}=0^{\circ}$), with (left) and without (right) applying the scaling procedure.
with $E$ being the shower primary energy. We find $a=57.4\,\rm g.cm^{-2}$, $c=421.9\,\rm g.cm^{-2}$ for proton-induced showers and $a=65.2\,\rm g.cm^{-2}$, $c=270.6\,\rm g.cm^{-2}$ for iron showers. From Eq 1, we can then derive the $X_{\rm max}$ coordinates of the target shower by numerically integrating the grammage along the shower axis, assuming Linsley’s atmospheric model. This $X_{\rm max}$ position is then assumed to be the point-like source of the target shower radio emission. We mention that we used here a parametrization of $X_{\rm max}>$ based on numerical simulations, instead of an empirical one, in order to get results as close as possible to ZHAireS simulations. The current Radio Morphing implements only proton and iron primaries, but it would be straightforward to implement other cosmic ray primaries using a similar parametrization.
@@ 3.4 Altitude at $X_{\mathrm{max}}$
Variations in the shower altitude also affect the air density at $X_{\rm max}$. The previous version of Radio Morphing was limited to a narrow zenith-angle range, being valid only for nearly horizontal showers ($\theta\gtrsim 86^{\circ}$). In the updated version, we account for the dependence of the radio emission on the air density at $X_{\rm max}$, which extends the range of validity to zenith angles between $60^{\circ}$ and $87.1^{\circ}$. Following [19], we model this dependency using a set of 11 00011\,000 ZHAireS simulations of cosmic-ray showers (with proton or iron primaries) on a star-shape layout, with zenith angles between $45^{\circ}$ and $90^{\circ}$, energies between $\rm 0.01\,EeV$ and $\rm 3.98\,EeV$ and 4 azimuth angles ($0^{\circ},90^{\circ},180^{\circ},270^{\circ}$). For each air shower, we reconstructed the charge excess and geomagnetic radiation energies $E_{\rm rad}^{\rm ce}$ and $E_{\rm rad}^{\rm geo}$, by separating each component and integrating their energy fluence following [20, 19]. These energies were then corrected from any dependency on the primary particle energy and on the geomagnetic angle using
@@
In Fig. 3, we show the geomagnetic and charge-excess emission dependency with air density at $X_{\rm max}$. The right-hand panel shows that the charge-excess amplitude increases almost linearly with the air density, as also shown in [20]. Indeed, a higher density provides more air atoms that can be ionized by the shower, leading to a larger build-up of negative charges, resulting in a stronger charge-excess emission. The left-hand panel shows that the geomagnetic emission first increases and then decreases when decreasing the air density at $X_{\rm max}$, i.e., for more inclined showers. When the shower inclination increases, charged particles can reach a larger drift velocity which initially results in a stronger transverse current and a larger geomagnetic emission [9]. Yet, if the shower is too inclined and the magnetic field is strong enough, the shower lateral extent at $X_{\rm max}$ becomes so large that the radio emission becomes incoherent, causing the drop observed for the lowest air densities as characterized in [21]. As a result, the charge-excess emission was fitted using a simple power law, while the geomagnetic emission was fitted using a broken power-law.
@@ 3.5 Shower-to-shower fluctuations
Due to the probabilistic nature of interactions occurring in air-showers, the radio-emission from various showers induced by primary particles with identical mass composition, energy, and arrival direction, presents some differences that we refer to as shower-to-shower fluctuations. These fluctuations are intrinsically modeled by Monte-Carlo simulations for which hadronic interactions are evaluated in a probabilistic way. However, Radio Morphing uses only a small sample of reference showers; hence, hadronic interactions are not recomputed when generating new simulations. This allows one to considerably reduce the computation time but requires further implementation of shower-to-shower fluctuations that are consistent with the simulation data. Another improvement of this new Radio Morphing version is the implementation of shower-to-shower fluctuations, which can now be enabled by the user. The modeling of these effects relies on parametrizations of: (1) $X_{\rm max}$ fluctuations and (2) the fluctuations of the number of particles in the shower.
@@ 3.6 Interpolation and extrapolation
The correction for the propagation effects assumes a point-link emission at $X_{\rm max}$, so that the radio emission electric field amplitude decreases linearly with the distance to $X_{\rm max}$. For fixed ($\omega$, $\phi_{\rm obs}$) angles, considering a scaled antenna at a distance $d_{\rm scaled}$ from $X_{\rm max}$ and a target antenna at a distance $d_{\rm target}$, our correction yields
@@ 4.1 Amplitude results
For each Radio Morphing test simulation, the radio signal was generated at 176 antenna positions located at an altitude of 1086 meters above sea level, and compared to the same outputs produced with ZHAireS simulations. We then quantified the efficiency of our method by evaluating the relative differences in amplitude of the electric field amplitude between ZHAireS and Radio Morphing simulations using
@@
where the brackets denote an average over time samples with amplitudes exceeding 20% of the ZHAireS peak value. This metric was preferred over a simple comparison of peak amplitudes, as it remains valid even for filtered signals, where the power is distributed over several samples. We also note that for each antenna, the dominant channel is used to evaluate the value of $\delta$. Using this metric, the mean value of $\delta$ over different signals quantifies the accuracy of our method (bias), while the standard deviation of $\delta$ evaluates the precision of Radio Morphing.
In the left-hand panel of Fig. 5, we show the mean (crosses) and standard deviation (blue shaded area) of the relative differences $\delta$ in the total peak amplitude at the antenna level, between the ZHAireS and Radio Morphing simulations, as a function of the target shower energy. We get mean relative differences close to $\delta=0\%$ and standard deviations of $\sim\delta=10\%$ for the most energetic showers, with energies above $1\,\rm EeV$, while both the precision and accuracy of Radio Morphing decrease for the lowest-energy showers. This loss of efficiency at low energy comes from two main effects. First, as mentioned in Section 2, our reference showers were run with a primary energy of $\mathcal{E}_{r}=3.98\,\rm EeV$, so the lowest primary energies in Fig. 5 are also the farthest from the reference one. Second, higher fluctuations of the radio emission are typically expected for low-energy showers in Monte Carlo simulations because of thinning effects.
The right-hand panel of Fig. 5 shows the distribution of $\delta$ for the three azimuth angles tested in our simulations. We observe a slight azimuth-dependent bias, as the distributions are not exactly centered on $\delta=0$, the largest deviations being observed for $\varphi=0^{\circ}$. This effect is still under investigation, but is probably related to the fact that showers with smaller geomagnetic angles, typically for $\varphi=0^{\circ}$ at the Dunhuang location, experience less coherence loss and therefore produce stronger radio emission after correction from the $\sin{\alpha}$ effect [21, 23]. However, our scaling of radio emission with the zenith angle shown in Fig. 3, which incorporates the coherence loss effects, relies on an average fit for all azimuth angles and therefore does not account for the azimuth-dependent effect. This will be corrected in further Radio Morphing implementations.
@@
Figure 8: Standard deviations of of the relative differences $\delta$ on the electric field peak amplitude, between ZHAireS and Radio Morphing simulations, as a function of the shower zenith angle. The results are shown for three different frequency bands: full band (black line), $50-200]\,\rm MHz$ (orange line) and $30-80]\,\rm MHz$ (blue line). The left-hand panel corresponds to a $3\sigma$ trigger threshold ($60\,\rm\mu V/m$), while the right-hand plot is for a $5\sigma$ trigger threshold ($110\,\rm\mu V/m$). The shaded area delimits the most inclined showers ($\theta>80^{\circ})$.
@@ 4.3 Timing
We also evaluated Radio Morphing performances on the peak-time reconstruction. A precise timing resolution is essential for the angular reconstruction of the arrival direction of the shower. For this, we evaluated the absolute peak-time difference, at the antenna level, between ZHAireS and Radio Morphing simulations. In the left-hand panel of Fig. 9, we show the distribution of the peak-time delay between Radio Morphing and ZHAireS simulations. For more than $99\%$ of antennas we get a time difference $\Delta t$ within $\pm 5{\rm ns}$, i.e., below the typical timing resolution of most radio experiments. This performance is allowed by the new interpolation method used in this version of Radio Morphing [12] and is a major improvement compared to the version of [19].

### v2

@@ Abstract
The preparation of the next-generation of large-scale radio experiments requires running a large number of simulations to explore multiple detector configurations over vast areas and develop novel methods for the reconstruction of air shower parameters. While Monte Carlo simulations are accurate and reliable tools, they are too computationally expensive to explore the full parameter space of these new detectors within a reasonable timescale, and faster and more efficient methods are needed. We introduce a new version of Radio Morphing, a semi-analytical tool designed to simulate the radio emission of any cosmic-ray induced air shower with zenith angle $\theta>60^{\circ}$, at any desired antenna position, from the simulation data of a few reference showers at given positions. This version incorporate refined scaling laws of the radio emission with the shower zenith angle, a novel interpolation method, the implementation of the charge-excess mechanism and the possibility to enable shower-to-shower fluctuations. We present the latest performances of Radio Morphing, tuned for a GRAND-like detector, which now provides an estimation of air shower radio signals four orders of magnitude faster than standard Monte Carlo simulations, while keeping an accuracy on the peak amplitude better than $17\%$ on unfiltered traces, better than $13\%$ in the [50 - 200] MHz band, and below $\sim 10\%$ in the [30 - 80] MHz band.
@@ 1 Introduction
Radio detection is now a mature, well-established technique, used by a wide range of experiments for cosmic ray or neutrino-induced air shower detection to unveil the high-energy Universe. Large-scale radio experiments such as GRAND [1], BEACON [2], SKA [3], or the radio upgrade of the Pierre Auger Observatory [4] are now being built to reach unprecedented sensitivity, field of view, and angular resolution probing the cosmic particle flux at the highest energies. This goal requires to run a large number of air-shower simulations to explore different antenna configurations, test reconstruction methods, and evaluate the detector performances. While Monte Carlo simulations such as ZHAireS [5, 6] and Coreas [7, 8] provide accurate simulation results, these are too computationally intensive to explore all the parameter space with large-scale experiments. On the other hand, macroscopic approaches such as MGMR [9, 10] are much faster but limited by our incomplete knowledge of the physical processes, particularly when looking for showers with very inclined arrival directions. This motivates the need to develop fast and accurate estimations of the radio signal. Already significant efforts have been made towards this direction, with template synthesis for vertical and mildly inclined showers [11, 12]. On the other hand, inclined showers remain challenging because of asymmetry effects and rapid changes in the emission with the shower zenith angle. We introduce here a new version of Radio Morphing, a semi-analytical tool designed to estimate the radio emission from any air shower with zenith angle above $60^{\circ}$, at any position, from the reference data of a few simulated showers at given positions. The current version incorporates shower-to-shower fluctuations, refined scaling laws, interpolation method, and reference library compared to the version of [13]. We present the current Radio Morphing method and its implementation in Sections 2 and 3 respectively, and show the latest Radio Morphing performances in Section 4.
@@ 2.1 First principles of Radio Morphing
The Radio Morphing procedure is performed on the full-band electric field traces, to infer general results that can then be filtered in the desired frequency range, similarly to Monte Carlo simulations.
@@ 2.2 Reference library and plane selection
When generating signals with Radio Morphing, the user must specify a target primary energy, zenith and azimuth angle $(E_{t},\,\theta_{t},\,\varphi_{t})$. We then choose, in our library, the reference ZHAireS simulation with the closest zenith angle $\theta_{\rm ref}$. The user must also specify a coordinate list of target antenna positions ($x_{t}$, $y_{t}$, $z_{t}$) in the AIRES reference frame, at which the radio signal will be computed. Hence, for each target antenna, we take within the selected ZHAireS simulation, the reference plane with $X_{\rm max}$ distance $D_{\rm plane}$ that is the closest to the $X_{\rm max}$-antenna distance $d_{\rm ant}$. This procedure is done to reduce as much as possible the error related to the extrapolation of the signal.
@@ 3.1 Primary composition
Figure 2: (Top) Lateral distribution functions (LDFs) for two showers with the same arrival directions ($\theta=75^{\circ},\,\varphi=90^{\circ}$), but different primary energies ($E_{r}=3.98\,\rm EeV,\,E_{t}=1.0\,\rm EeV$) without any energy scaling (top left) and with the energy scaling procedure (top right). The $\omega$ angle corresponds to the angular deviation from the shower axis, taken from $X_{\rm max}$ as shown in Fig. 1. $\Delta I/I$ gives the relative differences between the integral of both curves (see text). (Bottom) LDFs for two showers with the same primary energy and zenith angle ($E=3.98\,\rm EeV,\theta=75^{\circ}$), but different azimuth angles ($\varphi_{r}=90^{\circ},\,\varphi_{t}=0^{\circ}$), with (left) and without (right) applying the scaling procedure.
with $E$ being the shower primary energy. We find $a=57.4\,\rm g.cm^{-2}$, $c=421.9\,\rm g.cm^{-2}$ for proton-induced showers and $a=65.2\,\rm g.cm^{-2}$, $c=270.6\,\rm g.cm^{-2}$ for iron showers. From Eq (1), we can then derive the $X_{\rm max}$ coordinates of the target shower by numerically integrating the grammage along the shower axis, assuming Linsley’s atmospheric model. This $X_{\rm max}$ position is then assumed to be the point-like source of the target shower radio emission. We mention that we used here a parametrization of $X_{\rm max}>$ based on numerical simulations, instead of an empirical one, in order to get results as close as possible to ZHAireS simulations. The current Radio Morphing implements only proton and iron primaries, but it would be straightforward to implement other cosmic ray primaries using a similar parametrization.
@@ 3.4 Altitude at $X_{\mathrm{max}}$
Variations in the shower altitude also affect the air density at $X_{\rm max}$. The first version of Radio Morphing was limited to a narrow zenith-angle range, being valid only for nearly horizontal showers ($\theta\gtrsim 86^{\circ}$). In the updated version, we account for the dependence of the radio emission on the air density at $X_{\rm max}$, which extends the range of validity to zenith angles between $60^{\circ}$ and $87.1^{\circ}$. Following [20], we model this dependency using a set of 11 00011\,000 ZHAireS simulations of cosmic-ray showers (with proton or iron primaries) on a star-shape layout, with zenith angles between $45^{\circ}$ and $90^{\circ}$, energies between $\rm 0.01\,EeV$ and $\rm 3.98\,EeV$ and 3 azimuth angles ($0^{\circ},90^{\circ},180^{\circ}$). For each air shower, we reconstructed the charge excess and geomagnetic radiation energies $E_{\rm rad}^{\rm ce}$ and $E_{\rm rad}^{\rm geo}$, by separating each component and integrating their energy fluence following [21, 20]. These energies were then corrected from any dependency on the primary particle energy and on the geomagnetic angle using
@@
In Fig. 3, we show the geomagnetic and charge-excess emission dependency with air density at $X_{\rm max}$. The right-hand panel shows that the charge-excess amplitude increases almost linearly with the air density, as also shown in [21]. Indeed, a higher density provides more air atoms that can be ionized by the shower, leading to a larger build-up of negative charges, resulting in a stronger charge-excess emission. The left-hand panel shows that the geomagnetic emission first increases and then decreases when decreasing the air density at $X_{\rm max}$, i.e., for more inclined showers. When the shower inclination increases, charged particles can reach a larger drift velocity which initially results in a stronger transverse current and a larger geomagnetic emission [9]. Yet, if the shower is too inclined and the magnetic field is strong enough, the shower lateral extent at $X_{\rm max}$ becomes so large that the radio emission becomes incoherent, causing the drop observed for the lowest air densities as characterized in [22]. As a result, the charge-excess emission was fitted using a simple power law, while the geomagnetic emission was fitted using a broken power-law. These fits were done using the full band electric field, to get an average trend of the coherence loss effects.
@@ 3.5 Shower-to-shower fluctuations
Due to the probabilistic nature of interactions occurring in air-showers, the radio-emission from various showers induced by primary particles with identical mass composition, energy, and arrival direction, presents some differences that we refer to as shower-to-shower fluctuations. These fluctuations are intrinsically modeled by Monte-Carlo simulations for which hadronic interactions are evaluated in a probabilistic way. However, Radio Morphing uses only a small sample of reference showers; hence, hadronic interactions are not recomputed when estimating radio signals. This allows one to considerably reduce the computation time but requires further implementation of shower-to-shower fluctuations that are consistent with the simulation data. This new Radio Morphing version models these effects, at the first order, with parameterizations of $X_{\rm max}$ fluctuations and fluctuations in the number of particles in the shower. The user can now enable shower-to-shower fluctuations when generating radio signal estimations.
@@ 3.6 Interpolation and extrapolation
The correction for the propagation effects assumes a point-like emission at $X_{\rm max}$, so that the radio emission electric field amplitude decreases linearly with the distance to $X_{\rm max}$. For fixed ($\omega$, $\phi_{\rm obs}$) angles, considering a scaled antenna at a distance $d_{\rm scaled}$ from $X_{\rm max}$ and a target antenna at a distance $d_{\rm target}$, our correction yields
@@ 4.1 Amplitude results
For each Radio Morphing test simulation, the radio signal was generated at 176 antenna positions located at an altitude of 1086 meters above sea level corresponding to the location of Dunhuang (China), and compared to the same outputs produced with ZHAireS simulations. We then quantified the efficiency of our method by evaluating the relative differences in the electric field amplitude between ZHAireS and Radio Morphing simulations using
@@
where the brackets denote an average over time samples with amplitudes exceeding 20% of the ZHAireS peak value, considering a sampling rate of $0.5\,\rm ns$. This metric was preferred over a simple comparison of peak amplitudes, as it remains valid even for filtered signals, where the power is distributed over several samples. We also note that for each antenna, the dominant channel is used to evaluate the value of $\delta$. Using this metric, the mean value of $\delta$ over different signals quantifies the accuracy of our method (bias), while the standard deviation of $\delta$ evaluates the precision of Radio Morphing.
In the left-hand panel of Fig. 5, we show the mean (crosses) and standard deviation (blue shaded area) of the relative differences $\delta$ in the total peak amplitude at the antenna level, between the ZHAireS and Radio Morphing simulations, as a function of the target shower energy. We get mean relative differences close to $\delta=0\%$ and standard deviations of $\sim\delta=10\%$ for the most energetic showers, with energies above $1\,\rm EeV$, while both the precision and accuracy of Radio Morphing decrease for the lowest-energy showers. This loss of efficiency at low energy comes from two main effects. First, as mentioned in Section 2, our reference showers were run with a primary energy of $\mathcal{E}_{r}=3.98\,\rm EeV$, so the lowest primary energies in Fig. 5 are also the farthest from the reference one. Second, at low energy the cascade involves fewer particle interactions, resulting in reduced statistical averaging of fluctuations and therefore a larger intrinsic dispersion in the radio signal and in the relative differences $\delta$.
The right-hand panel of Fig. 5 shows the distribution of $\delta$ for the three azimuth angles tested in our simulations, restricting to showers with primary energy above $2\,\rm EeV$. We observe a slight azimuth-dependent bias, as the distributions are not exactly centered on $\delta=0$, the largest deviations being observed for $\varphi=0^{\circ}$. This effect is still under investigation. We hypothesize that the scaling of the radio emission with zenith angle shown in Fig. 3 is in fact azimuth-dependent. Showers with smaller geomagnetic angles — for instance $\varphi=0^{\circ}$ at the Dunhuang site — experience reduced coherence losses and therefore yield stronger radio emission than showers with larger geomagnetic angles [22, 23]. The current zenith-angle scaling implemented in Radio Morphing, which accounts for coherence loss effects, effectively represents an average over all azimuth angles. This limitation can be addressed by extending the reference library to include showers covering a range of azimuth angles, or by introducing an azimuth-dependent zenith-angle scaling. Such improvements will be implemented in future developments of the method.
@@
Figure 8: Standard deviations of of the relative differences $\delta$ on the electric field peak amplitude, between ZHAireS and Radio Morphing simulations, as a function of the shower zenith angle. The results are shown for three different frequency bands: full band (black line), $50-200]\,\rm MHz$ (orange line) and $30-80]\,\rm MHz$ (blue line). The left-hand panel corresponds to a $3\sigma$ trigger threshold ($60\,\rm\mu V/m$), while the right-hand plot is for a $5\sigma$ trigger threshold ($110\,\rm\mu V/m$). The shaded area delimits the most inclined showers ($\theta>80^{\circ})$. The top panels show the differences in zenith angle $\delta\theta$ between the reference shower used and the target shower for which the radio signal is generated.
@@ 4.3 Timing
We also evaluated Radio Morphing performances on the peak-time reconstruction. A precise timing resolution is essential for the angular reconstruction of the arrival direction of the shower. For this, we evaluated the absolute peak-time difference, at the antenna level, between ZHAireS and Radio Morphing simulations. In the left-hand panel of Fig. 9, we show the distribution of the peak-time delay between Radio Morphing and ZHAireS simulations. Radio Morphing outputs are generated with a default sampling rate of $0.5\,\rm ns$. For more than $99\%$ of antennas we get a time difference $\Delta t$ within $\pm 5{\rm ns}$, i.e., below the typical timing resolution of most radio experiments. This performance is allowed by the new interpolation method used in this version of Radio Morphing [13] and is a major improvement compared to the version of [20].